OpenStudy (anonymous):

Part 1: Explain, using complete sentences, how to simplify the expression below. Part 2: What is the simplified expression? (3x3y4)(2x2y6)

jhonyy9 (jhonyy9):

3*2=6 x3*x2=x(3+2)=x5 y4*y6=y(4+6)=y10

OpenStudy (hoblos):

multiply the coefficient and add the powers of x and y you will get 6x5y10